OpenStudy (anonymous):

Solve 10p-2(3p-6)=4(5p-6)-10p in steps?

OpenStudy (pfenn1):

\[10p-2(3p-6) =4(5p-6)-10p\]\[10p-6p+12=20p-24-10p\]\[4p+12=10p-24\]\[6p=36\]\[p=6\]

OpenStudy (anonymous):

thank you.

OpenStudy (pfenn1):

you're welcome